A ball is thrown horizontally from the roof of a 20.0 m high building and lands 10.0 m away from the base of the building. Determine initial and final velocity.


1
Expert's answer
2021-02-05T02:13:52-0500

a) Let's first find the time that the ball takes to reach the ground:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot20\ m}{9.8\ \dfrac{m}{s^2}}}=2.0\ s"

Then, the initial horizontal velocity can be found as follows:


"v_{0x}=\dfrac{x}{t}=\dfrac{10\ m}{2.0\ s}=5\ \dfrac{m}{s}."

b) The initial horizontal velocity remains unchainged during the flight of the ball. Let's find the final vertical velocity of the ball:


"v_y=v_{0y}-gt=0-9.8\ \dfrac{m}{s^2}\cdot2.0\ s=-19.6\ \dfrac{m}{s}."

Then, we can find the final velocity of the ball from the Pythagorean theorem:


"v=\sqrt{v_x^2+v_y^2}=\sqrt{(5\ \dfrac{m}{s})^2+(-19.6\ \dfrac{m}{s})^2}=20.23\ \dfrac{m}{s}."